OpenStudy (anonymous):

Please answer. Convert the polar equation θ = 3 π/4 to a Cartesian equation. a. y= x b. y= -x c. y= 3π/4x d. y= 3π/4

OpenStudy (anonymous):

remember that tan(theta) = y/x this meants that y = tan(theta)x in our case y= tan(3pi/4)x